# Reduction of Cosymplectic groupoids by cosymplectic moment maps

### v2

@@ 4.2. Reduction of the symplectic leaf
[…]
commutes. The reduced symplectic groupoid $\mathcal{J}_{0}^{-1}(0)/G$ is constructed by considering the symplectic structure as in [23], which is compatible with the structural maps of the groupoid [17, Prop. 4.6]. As a direct consequence we get the next theorem.
@@
Let $G$ be a Lie group acting on a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$, and $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ a cosymplectic moment map satisfying the conditions of Theorem 3.4. Let $\Sigma$ the symplectic leaf containing $M$. Then
@@
Since $G$ acts freely and properly on $M$, the $G$-action on $\Sigma$ is also free and proper [17]. Thus, by a similar argument as in Remark 2.5, 0 is a regular value of $\mathcal{J}_{0}$. Moreover, $(\mathcal{J}_{0}^{-1}(0)/G,(L^{*}\omega)_{red})\rightrightarrows M/G$ is a symplectic groupoid [17]. The characterization of the reduced symplectic form on $\mathcal{J}_{0}^{-1}(0)/G$ is
@@
where $L_{0}:\mathcal{J}_{0}^{-1}(0)\to\Sigma$ is the inclusion map and $P_{0}:\mathcal{J}_{0}^{-1}(0)\to\mathcal{J}_{0}^{-1}(0)/G$ is the quotient map. By the commutative in (11), we have
@@ 4.3. Reduction of the multiplicative Chern class
In [15] the authors establish a relation between $S^{1}$-central extensions with vanishing multiplicative Chern class and cosymplectic groupoids. We use this identification and Theorem 3.4 to show that if the multiplicative Chern class vanishes, the same happens in the quotient.
Let $(\mathcal{G},\omega)\rightrightarrows M$ be a Lie groupoid with a closed multiplicative 2-form, such that for any $x\in M$, $\text{rank}(\omega_{x})=2\dim(M)$. The pair $(\mathcal{G},\omega)$ is known as an oversymplectic groupoid [15, 7]. Furthermore, we suppose that $\mathcal{G}$ is orientable and proper. Assume that the foliation defined by $\ker\omega$ is simple, i.e., the leaf space $\mathcal{G}/\ker\omega$ is a smooth manifold and the map $\mathcal{G}\to\mathcal{G}/\ker\omega$ is a surjective submersion. Furthermore, assume that the corank of the foliation $\ker\omega$ is 1. Under these conditions, there exists a $S^{1}$-central extension
[…]
where $(\Sigma_{1},\omega_{1})$ is a symplectic groupoid and $q^{*}\omega_{1}=\omega$.
Moreover, $\mathcal{G}$ becomes a $S^{1}$-principal bundle. Let $\partial_{\theta}$ be the infinitesimal generator of the $S^{1}$- action on $\mathcal{G}$. Because $\mathcal{G}$ is a proper groupoid there exists a multiplicative form $\eta\in\Omega^{1}(\mathcal{G})$ such that $\imath_{\partial_{\theta}}\eta=1$. The form $\eta$ is related to the multiplicative Ehresmann connection, and the curvature of this connection is $\Omega=d\eta$. For more details on this construction the reader can consult [15, 16].
The 2-form $\Omega$ is multiplicative and $q$-basic i.e., there exists a form $\Omega_{1}\in\Omega_{cl}(\Sigma_{1})$ such that $q^{*}\Omega_{1}=\Omega$. The multiplicative Chern class is defined as the cohomology class associated with $\Omega_{1}$ (in the multiplicative de Rham cohomology). One of the main results in [15, Thm. 4.12] states that there exists a multiplicative 1-form $\eta$ such that $(\mathcal{G},\omega,\eta)$ is a cosymplectic groupoid if and only if the corresponding $S^{1}$-central extension has vanishing multiplicative Chern class.
For a cosymplectic groupoid the foliation defined by the kernel of the 2-form is simple. On the other hand, if $\mathcal{G}$ is a proper groupoid, and there is an $G$-action on $\mathcal{G}$ and a cosymplectic moment map $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ as in Theorem 3.4, the groupoid $\mathcal{G}_{red}\rightrightarrows M/G$ may be a non-proper groupoid. However, with additional conditions the reduced groupoid is proper.
@@
[…]
@@
Since the action $G$ on $\mathcal{H}$ is free and proper [17], the quotients $\mathcal{H}/G$ and $M/G$ are smooth manifolds. Provided that the structural maps of $\mathcal{H}\rightrightarrows M$ are $G$-equivariant, then they descend to $\mathcal{H}/G$. Now, we show that under these hypotheses $\mathcal{H}$ is a proper groupoid. To show that $(s,t):\mathcal{H}\to M\times M$ is a proper map, we consider a sequence of arrows $y_{n}\xleftarrow{g_{n}}x_{n}\in\mathcal{H}$ such that $(x_{n},y_{n})\to(x,y)\in M\times M$, and we show that $g_{n}$ admits a convergent subsequence in $\mathcal{H}$ (see [12, 4]). Since $\mathcal{G}$ is a proper groupoid, there exists a subsequence $g_{n_{k}}\to g\in\mathcal{G}$. Because $\mathcal{H}\subset\mathcal{G}$ is closed, then $g\in\mathcal{H}$.
Let $[y_{n}]\xleftarrow{[g_{n}]}[x_{n}]\in\mathcal{H}/G$ such that $([x_{n}],[y_{n}])\to([x],[y])\in M/G\times M/G$. Thus, it is enough to show that $[g_{n}]$ admits a convergent subsequence in $\mathcal{H}/G$. Since $(p,p):M\times M\to M/G\times M/G$ and $(s,t):\mathcal{H}\to M\times M$ are proper maps, there exists a subsequence $y_{n}\xleftarrow{g_{n_{k}}}x_{n}$ converging to $g\in\mathcal{H}$. By the continuity of the map $\mathcal{H}\to\mathcal{H}/G$, we conclude that $[g_{n_{k}}]\to[g]\in\mathcal{H}/G$. ∎
In Lemma 4.9, considering the group $G$ to be compact, the condition that the map $M\to M/G$ be proper is changed to be a closed map. Next, we state a result which follows directly from Theorem 3.4 and the characterization between cosymplectic groupoids and $S^{1}$-central extension with vanishing multiplicative Chern class presented in [15, Thm. 4.12].
@@
Let $G$ be a Lie group acting on a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$, and $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ a cosymplectic moment map satisfying the conditions of Theorem 3.4. If $p:M\to M/G$ is a proper map, then for $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$ the corresponding $S^{1}$-central extension has vanishing multiplicative Chern class.
@@
By using Theorem 3.4 we obtain a cosymplectic groupoid $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$. Since $M\to M/G$ was assumed to be a proper map, by Lemma 4.9, $\mathcal{G}_{red}$ is proper. Consequently, by using the result in [15, Thm. 4.12], we conclude that the corresponding $S^{1}$-central extension of $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$ has vanishing multiplicative Chern class. ∎
@@
Let $(\mathcal{G},\omega)\rightrightarrows M$ be a corank 1, orientable, proper oversymplectic groupoid, with simple foliation $\ker\omega$, such that $H^{1}(\mathcal{G})=0$. Let $G$ be a compact Lie group acting on $\mathcal{G}$ by Lie groupoid automorphisms preserving $\omega$. The following is true:
@@
    If the corresponding $S^{1}$-central extension has vanishing multiplicative Chern class, then there exists a cosymplectic moment map $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ for a cosymplectic groupoid structure on $\mathcal{G}$, such that $\xi(\langle\mathcal{J},v\rangle)=0$ for all $v\in\mathfrak{g}$.
@@
    If, in addition, $\mathcal{J}$ is a groupoid morphism, $G$ acts freely on $M$, and $M\to M/G$ is a closed map, then for $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$ the corresponding $S^{1}$-central extension has vanishing multiplicative Chern class.
@@
To prove item (1), the condition on the multiplicative Chern class defines a cosymplectic structure $(\mathcal{G},\omega,\eta)$ [15, Thm. 4.12]. Because $G$ is compact, we consider the average
@@
where $d\Lambda$ is the Haar measure of the compact group $G$. From Haar measure invariance it follows that $\bar{\eta}$ is a $G$-invariant, closed, multiplicative 1-form in $\mathcal{G}$.
We claim that $(\mathcal{G},\omega,\bar{\eta})$ is a cosymplectic manifold. In fact, note that the Reeb vector field $\xi$ of $(\mathcal{G},\omega,\eta)$ satisfies $\imath_{\xi}\omega=0$ and $\imath_{\xi}\bar{\eta}\neq 0$. Thus, if the map $\flat:\mathfrak{X}(\mathcal{G})\to\Omega^{1}(\mathcal{G})$ defined in (1) is not injective, we can consider $X\neq 0$ in the kernel of $\flat$. Note that $0=\imath_{\xi}\flat(X)=(\imath_{X}\bar{\eta})(\imath_{\xi}\bar{\eta})$. Hence, $\bar{\eta}$ vanishes on $X$, and $\imath_{X}\omega=\flat(X)=0$. Thus, $X=f\xi$ for some non-vanishing smooth function $f$. However, $\imath_{X}\bar{\eta}=f\imath_{\xi}\bar{\eta}$, which is a contradiction. Consequently, $\flat$ is an isomorphism. We conclude this claim, by considering the normalization $\bar{\xi}$ of $\xi$, such that $\imath_{\bar{\xi}}\bar{\eta}=1$, and Remark 2.2.
Since $\bar{\eta}$ is invariant by the action, we get that $d(\imath_{v_{\mathcal{G}}}\bar{\eta})=\mathcal{L}_{v_{\mathcal{G}}}\bar{\eta}=0$. Because $\epsilon^{*}\bar{\eta}=0$, then $0=\imath_{v_{M}}\epsilon^{*}\bar{\eta}|_{x}=\imath_{v_{\mathcal{G}}}\bar{\eta}|_{\epsilon(x)}$ for $x\in M$. Hence, $\imath_{v_{\mathcal{G}}}\bar{\eta}=0$ for all $v\in\mathfrak{g}$. The existence of the cosymplectic moment map $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ follows form Proposition 2.6.
To prove item (2), we consider the reduced cosymplectic groupoid $(\mathcal{G}_{red},\omega_{red},\bar{\eta}_{red})\rightrightarrows M/G$ obtained from Theorem 3.4. By using Lemma 4.9, $\mathcal{G}_{red}$ is proper. From [15, Thm. 4.12], we conclude that the corresponding $S^{1}$-central extension of $(\mathcal{G}_{red},\omega_{red},\bar{\eta}_{red})\rightrightarrows M/G$ has vanishing multiplicative Chern class. ∎
5. Further direction of study
  • Recall that in the cosymplectic reduction Theorem 2.4 it is assumed that $v_{Q}$ is the Hamiltonian vector field of $\langle J,v\rangle$ and $\xi\langle J,v\rangle=0$, this in particular implies that $\mathcal{L}_{v_{Q}}\omega=0$ and $\mathcal{L}_{v_{Q}}\eta=0$. However, we can suppose conformal symmetries, which leads us to consider the conditions
    for some non-vanishing function $f_{v}$ for each $v\in\mathfrak{g}$. In this context, the reduction theorems are different, and would state some other relation with respect to the cosymplectic groupoids reduction.
@@ 5. Further direction of study
  • In [15, Prop. 3.3] the authors state an infinitesimal identification with a Lie bialgebroid of the form $(T^{*}M\oplus\mathbb{R},TM\oplus\mathbb{R})$. However it is also known the existence of generalized Lie bialgebroids [21] and triangular generalized Lie bialgebroids as in [20]. The reduction results of this paper can be used to reduce such bialgebroids [21]. They can also be used to compare other geometric terms, such as the modular class in the reduction presented in [20].
@@
  • Integration of Poisson structures can also be understand via Poisson sigma models [10]. As we mention previously, if the groupoid $\mathcal{G}$ is cosymplectic, it has a Poisson bracket. Furthermore, the units $M$ also has a Poisson structure and a Lie algebroid structure. Results in [15] and in these notes can be extended to the integration theory given in [10] via sigma models.
@@
  • The construction of the moment map in Section 4.1 comes from the ideas in [6], where there is an action of a Lie group on an algebroid which lifts an action on the basis. By the Lie’s second theorem, this action is integrated to an action of a Lie group on a Lie grouproid. It would be valuable to see if the ideas worked out here can be generalized by considering the action of a 2-group on a cosymplectic groupoid and its infinitesimal counterpart.

### v3

@@ 4.2. Reduction of the symplectic leaf
${\mathcal{J}_{0}^{-1}(0)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{L}$$\scriptstyle{L_{0}}$$\scriptstyle{P_{0}}$${\mathcal{J}^{-1}(0)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$$\scriptstyle{P}$${\mathcal{J}_{0}^{-1}(0)/G\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{L_{red}\hskip 14.22636pt}$${\mathcal{G}_{red}}$${(\Sigma,L^{*}\omega)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{L}$${(\mathcal{G},\omega,\eta).}$
commutes. The reduced symplectic groupoid $\mathcal{J}_{0}^{-1}(0)/G$ is constructed by considering the symplectic structure as in [MW], which is compatible with the structural maps of the groupoid [FOR, Prop. 4.6]. As a direct consequence we get the next theorem.
@@
Let $G$ be a Lie group acting on a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$ by cosymplectic groupoid maps, and let $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ be a cosymplectic moment map satisfying the conditions of Theorem 3.4. Let $\Sigma$ the symplectic leaf containing $M$. Then
@@
Since $G$ acts freely and properly on $M$, the $G$-action on $\Sigma$ is also free and proper [FOR]. Thus, by a similar argument as in Remark 2.5, 0 is a regular value of $\mathcal{J}_{0}$. Moreover, $(\mathcal{J}_{0}^{-1}(0)/G,(L^{*}\omega)_{red})\rightrightarrows M/G$ is a symplectic groupoid [FOR]. The reduced symplectic form $(L^{*}\omega)_{red}$ on $\mathcal{J}_{0}^{-1}(0)/G$ is characterized by
@@
where $L_{0}:\mathcal{J}_{0}^{-1}(0)\to\Sigma$ is the inclusion map and $P_{0}:\mathcal{J}_{0}^{-1}(0)\to\mathcal{J}_{0}^{-1}(0)/G$ is the quotient map. By the commutativity in (12), we have
@@ 4.3. Reduction of the multiplicative Chern class
As mentioned in Section 2.2.2, [FI] establishes a relation between $S^{1}$-central extensions with vanishing multiplicative Chern class and cosymplectic groupoids. We use this identification and Theorem 3.4 to show that if the multiplicative Chern class vanishes, the same happens in the quotient.
Let $(\mathcal{G},\omega)$ be a corank 1, orientable, proper oversymplectic groupoid over $M$, whose foliation $\ker\omega$ is simple. As in Theorem 2.11, there is a one-to-one correspondence between a cosymplectic groupoid structure on $(\mathcal{G},\omega)$ and the vanishing of the multiplicative Chern class. Thus, if $\mathcal{G}$ is a cosymplectic groupoid, we are interested in understanding the behavior of the multiplicative Chern class under the reduction procedure described in Theorem 3. Assume a $G$-action on $\mathcal{G}$ and a cosymplectic moment map $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ as in Theorem 3.4. Note that although $\mathcal{G}$ is proper, the reduced groupoid $\mathcal{G}_{red}\rightrightarrows M/G$ may be a non-proper groupoid. However, with additional conditions the reduced groupoid is proper.
@@
Let $G$ be a Lie group acting on a proper Lie groupoid $\mathcal{G}\rightrightarrows M$ by groupoid morphisms. Let $\mathcal{H}\rightrightarrows M$ be a closed Lie subgroupoid of $\mathcal{G}$. Suppose that $\mathcal{H}$ is closed under the action of $G$. If $G$ acts freely and properly on $M$, and the map $M\xrightarrow{p}M/G$ is proper, then $\mathcal{H}/G\rightrightarrows M/G$ is a proper Lie groupoid.
@@
Since the action of $G$ on $\mathcal{H}$ is free and proper [FOR], the quotients $\mathcal{H}/G$ and $M/G$ are smooth manifolds. As the structural maps of $\mathcal{H}\rightrightarrows M$ are $G$-equivariant, then they descend to $\mathcal{H}/G$. Now, we show that under these hypotheses $\mathcal{H}$ is a proper groupoid. To show that $(s,t):\mathcal{H}\to M\times M$ is a proper map, we consider a sequence of arrows $y_{n}\xleftarrow{g_{n}}x_{n}\in\mathcal{H}$ such that $(x_{n},y_{n})\to(x,y)\in M\times M$, and we show that $g_{n}$ admits a convergent subsequence in $\mathcal{H}$ (see [dH, Bour]). Since $\mathcal{G}$ is a proper groupoid, there exists a subsequence $g_{n_{k}}\to g\in\mathcal{G}$. Because $\mathcal{H}\subset\mathcal{G}$ is closed, then $g\in\mathcal{H}$.
Let $y_{n}]\xleftarrow{[g_{n}]}[x_{n}]\in\mathcal{H}/G$ be such that $([x_{n}],[y_{n}])\to([x],[y])\in M/G\times M/G$. We want to show that $g_{n}]$ admits a convergent subsequence in $\mathcal{H}/G$. Since $(p,p):M\times M\to M/G\times M/G$ and $(s,t):\mathcal{H}\to M\times M$ are proper maps, there exists a subsequence $y_{n}\xleftarrow{g_{n_{k}}}x_{n}$ converging to $g\in\mathcal{H}$. By the continuity of the map $\mathcal{H}\to\mathcal{H}/G$, we conclude that $g_{n_{k}}]\to[g]\in\mathcal{H}/G$. ∎
Next, we state a result which follows directly from Theorem 3.4 and the characterization of cosymplectic groupoids in terms of $S^{1}$-central extensions with vanishing multiplicative Chern class presented in Theorem 2.11.
@@
Let $G$ be a Lie group acting on a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$, and let $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ be a cosymplectic moment map satisfying the conditions of Theorem 3.4. If $p:M\to M/G$ is a proper map, then for $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$ the corresponding $S^{1}$-central extension has vanishing multiplicative Chern class.
@@
By using Theorem 3.4 we obtain a cosymplectic groupoid $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$. Since $M\to M/G$ was assumed to be a proper map, by Lemma 4.9, $\mathcal{G}_{red}$ is proper. Consequently, by using Theorem 2.11, we conclude that the corresponding $S^{1}$-central extension of $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$ has vanishing multiplicative Chern class. ∎
@@
Let $(\mathcal{G},\omega)\rightrightarrows M$ be a corank 1, orientable, proper oversymplectic groupoid, with simple foliation $\ker\omega$, such that $H^{1}(\mathcal{G})=0$. Let $G$ be a compact, connected Lie group acting on $\mathcal{G}$ by Lie groupoid automorphisms preserving $\omega$. The following is true:
@@
    If the corresponding $S^{1}$-central extension has vanishing multiplicative Chern class, then there exists a cosymplectic groupoid $(\mathcal{G},\omega,\bar{\eta})$ and a moment map $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$, such that $\xi(\langle\mathcal{J},v\rangle)=0$ for all $v\in\mathfrak{g}$.
@@
    If, in addition, $\mathcal{J}$ is a groupoid morphism, $G$ acts freely on $M$, and $M\to M/G$ is a closed map, then for $(\mathcal{G}_{red},\omega_{red},\bar{\eta}_{red})\rightrightarrows M/G$ the corresponding $S^{1}$-central extension has vanishing multiplicative Chern class.
@@
To prove item (1), we find an $G$-invariant multiplicative 1-form $\bar{\eta}$ on $\mathcal{G}$, such that $(\mathcal{G},\omega,\bar{\eta})$ is a cosymplectic groupoid. Namely, the condition on the multiplicative Chern class defines a cosymplectic structure $(\mathcal{G},\omega,\eta)$ (cf. Theorem 2.11). Because $G$ is compact, we consider the average
$\bar{\eta}=\int_{G}(\varphi^{\mathcal{G}}_{g})^{*}\eta d\Lambda_{g},$
where $d\Lambda$ is the Haar measure of the compact group $G$. From Haar measure invariance it follows that $\bar{\eta}$ is a $G$-invariant, closed, multiplicative 1-form in $\mathcal{G}$. Let $\xi$ be the Reeb vector field of $(\mathcal{G},\omega,\eta)$. Note that $d\varphi^{\mathcal{G}}_{g}\xi=f_{g}\xi$ for some non-vanishing $f_{g}\in C^{\infty}(\mathcal{G})$, thus $\imath_{\xi}\bar{\eta}=\int_{G}f_{g}d\Lambda_{g}$. Furthermore, if $f_{g}(h)=0$ for some $h\in\mathcal{G}$, because $\varphi^{\mathcal{G}}_{g}$ is a diffeomorphism, we conclude that $\xi(h)=0$, which cannot happens. Consequently, by the connectedness of $G$, $f_{g}$ has the same sign for all $g\in G$.
We claim that $(\mathcal{G},\omega,\bar{\eta})$ is a cosymplectic manifold. In fact, $\xi$ satisfies $\imath_{\xi}\omega=0$ and $\imath_{\xi}\bar{\eta}\neq 0$. Thus, if the map $\flat:\mathfrak{X}(\mathcal{G})\to\Omega^{1}(\mathcal{G})$ defined in (1) associated with $\bar{\eta}$ is not injective, then we can consider $X\neq 0$ in the kernel of $\flat$. Note that $0=\imath_{\xi}\flat(X)=(\imath_{X}\bar{\eta})(\imath_{\xi}\bar{\eta})$. Hence, $\bar{\eta}$ vanishes on $X$, and $\imath_{X}\omega=\flat(X)=0$. Thus, $X=f\xi$ for some non-vanishing smooth function $f$. However, $\imath_{X}\bar{\eta}=f\imath_{\xi}\bar{\eta}$, which is a contradiction. Consequently, $\flat$ is an isomorphism. We conclude this claim, by considering the normalization $\bar{\xi}$ of $\xi$, such that $\imath_{\bar{\xi}}\bar{\eta}=1$, and Remark 2.1.
Since $\bar{\eta}$ is invariant by the action, we get that $d(\imath_{v_{\mathcal{G}}}\bar{\eta})=\mathcal{L}_{v_{\mathcal{G}}}\bar{\eta}=0$. Because $\epsilon^{*}\bar{\eta}=0$, then $0=\imath_{v_{M}}\epsilon^{*}\bar{\eta}|_{x}=\imath_{v_{\mathcal{G}}}\bar{\eta}|_{\epsilon(x)}$ for $x\in M$. Hence, $\imath_{v_{\mathcal{G}}}\bar{\eta}=0$ for all $v\in\mathfrak{g}$. The existence of the cosymplectic moment map $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ follows from Proposition 2.6.
To prove item (2), we consider the reduced cosymplectic groupoid $(\mathcal{G}_{red},\omega_{red},\bar{\eta}_{red})\rightrightarrows M/G$ obtained from Theorem 3.4. In Lemma 4.9, considering that the group $G$ is compact, the condition that the function $M\to M/G$ is proper is replaced by that of being a closed map, therefore $\mathcal{G}_{red}$ is proper. From Theorem 2.11, we conclude that the corresponding $S^{1}$-central extension of $(\mathcal{G}_{red},\omega_{red},\bar{\eta}_{red})\rightrightarrows M/G$ has vanishing multiplicative Chern class. ∎
5. Further direction of study
  1. (1)
    In the cosymplectic reduction Theorem 2.4 it is assumed that $v_{Q}$ is the Hamiltonian vector field of $\langle J,v\rangle$ and $\xi\langle J,v\rangle=0$, this in particular implies that $\mathcal{L}_{v_{Q}}\omega=0$ and $\mathcal{L}_{v_{Q}}\eta=0$. However, we can suppose conformal symmetries, which leads us to consider the conditions
    for some non-vanishing function $f_{v}$ for each $v\in\mathfrak{g}$. In this context, the reduction theorems are different, and would state others relation with respect to the cosymplectic groupoids reduction. More generally, we may consider two broader interpretations of conformal symmetries in cosymplectic structures, as we present below.
    • A cosymplectic structure can be regarded as a symplectic structure on a line bundle $L\to M$ (cf. [TVY]). It is given by a 2-form $\omega\in\Omega^{2}(DL,\mathbb{R}_{M})$, where $DL$ denotes the Atiyah algebroid and $\mathbb{R}_{M}$ is the trivial line bundle. In this setting, a symplectic structure on $M$ with values in $L$ is a pair $(L,\omega)$, where $\omega$ is non-degenerate and closed with respect to the trivial representation of $DL$ on $\mathbb{R}_{M}$.
      In this context, one expects a cosymplectic groupoid to be a line VB-groupoid
@@ 5. Further direction of study
      together with a multiplicative cosymplectic structure $\omega\in\Omega^{2}(DL,\mathbb{R}_{\mathcal{G}})$. The approach to conformal symmetries in this framework is motivated by the trivial line bundle $L=\mathbb{R}\times M$, where a cosymplectic isomorphism with values in $L\to M$ can be translated into a diffeomorphism $\phi:M\to M$ together with a smooth function $g$ such that
      $\omega=\phi^{*}\omega+dg\wedge\phi^{*}\eta,\qquad\eta=\phi^{*}\eta.$
      In this way, the action of a Lie group leads to
      $L_{u_{M}}\omega=dG_{u}\wedge\eta,\qquad L_{u_{M}}\eta=0,$
      where the functions $G_{u}$ provide a natural candidate for a moment map of the action.
    • A cosymplectic structure can be regarded as a closed 2-form that restricts to a symplectic form on an involutive corank-one distribution $H\subset TM$. [^4] Contact geometry appears as the non-integrable case. Multiplicativity should then be interpreted as a condition on the distribution $H\subset T\mathcal{G}$, requiring it to define an involutive wide subgroupoid (cf. [CSa]). A cosymplectic isomorphism should preserve the distribution; that is, it consists of a diffeomorphism $\phi$ together with a smooth function $g$ such that
      $\omega=\phi^{*}\omega,\qquad\text{and}\qquad\phi^{*}\eta=g\,\eta,$
      since we must have $\phi_{*}H\subseteq H$. For a Lie group action, this translates into
      $L_{u_{M}}\omega=0,\qquad L_{u_{M}}\eta=G_{u}\,\eta.$
  2. (2)
@@
    In [FI, Prop. 3.3], given an IMcs-algebroid $(A,\mu,\nu,\zeta)$, the authors establish an isomorphism of Lie bialgebroids $(A,A^{*})\simeq(T^{*}M\oplus\mathbb{R},TM\oplus\mathbb{R}).$ However it is also known the existence of generalized Lie bialgebroids [IM] and triangular generalized Lie bialgebroids as in [IMLP]. The reduction results of this article can be used to reduce such bialgebroids. They can also be used to compare other geometric terms, such as the modular class in the reduction in [IMLP].
  3. (3)
@@
    Integration of Poisson structures can also be understand via Poisson sigma models [CF]. If the groupoid $\mathcal{G}$ is cosymplectic, it has a Poisson bracket. Furthermore, the units $M$ also has a Poisson structure and a Lie algebroid structure. Results in [FI] and in these notes can be extended to the integration theory given in [CF] via sigma models.
  4. (4)
@@
    The construction of the moment map in Section 4.1 comes from the ideas in [BC2], where there is an action of a Lie group on an algebroid which lifts an action on the basis. By the Lie’s second theorem, this action is integrated to an action of a Lie group on a Lie grouproid. It would be valuable to see if the ideas worked out here can be generalized by considering the action of a 2-group on a cosymplectic groupoid and its infinitesimal counterpart.
@@
Daniel López Garcia
Universidade de Federal Fluminense (UFF),
São Domingos, Niterói, 24210-200, Rio de Janeiro, Brazil.
danielflg@id.uff.br
Orcid: 0000-0002-8436-2599
Nicolas Martinez Alba
Universidad Nacional de Colombia (UNAL),
Av- Cra 30 #45-3, Bogotá 16486, Bogotá, Colombia.
nmartineza@unal.edu.co
Orcid: 0000-0002-2697-3228